If log4(y+2)=3, what is the value of y?
Answer by AnlytcPhil(1116)

If log4(y + 2) = 3, what is the value of y?

You need to learn this principle of logs:

To change a log equation into an exponential equation:

logBA = C can be replaced by the equivalent equation BC = A

By that principle, the equation

log4(y + 2) = 3 can be replaced by the equivalent equation:

43 = y + 2

64 = y + 2

62 = y 

Edwin
AnlytcPhil@aol.com